### v2

# Appendix for FANToM: A Framework to Align Neural Topic Models

[…]
ijcai24
Appendix for FANToM: A Framework to Align Neural Topic Models
[…]
Appendix A Modeling Details
Let $x$ be the BoW representation of a document. Let $V$ and $K$ be vocabulary size and no of topics respectively. The common encoder used across different models is:
Encoder:
$\begin{split}X_{1}&=\text{ReLU}(L^{1}\cdot x+L_{\text{bias}}^{1})\\ X_{2}&=\text{dropout}(X_{1},p_{\text{keep}})\\ X_{3}&=\text{BatchNorm}(L^{2}\cdot X_{2}+L_{\text{bias}}^{2})\\ \text{enc\_out}&=\text{Softplus}(X_{3}),\end{split}$
A common linear decoder with Dirichlet prior is given as:
Decoder:
$\begin{split}X_{4}&=\text{BatchNorm}(L^{3}\cdot z+L_{\text{bias}}^{3})\\ \text{recon}&=\text{LogSoftmax}(X_{4})\end{split}$
Certain models (such as ETM) use an embedding decoder. $\alpha$ and $\delta$ represent topic and word embeddings, respectively:
Decoder ETM:
$\begin{split}X_{4}&=z\cdot\eta\cdot\delta\\ \text{recon}&=\text{LogSoftmax}(\text{BatchNorm}(X_{4}))\end{split}$
For FANToM we have two decoders for text and authors of the document.
Decoder FANToM:
$\begin{split}X_{4}&=\text{BatchNorm}(L^{3}\cdot z+L_{\text{bias}}^{3})\\ X_{5}&=\text{BatchNorm}(L^{4}\cdot z+L_{bias}^{4})\\ \text{recon\_doc}&=\text{LogSoftmax}(X_{4})\\ \text{recon\_author}&=\text{LogSoftmax}(X_{5})\end{split}$
The hyperparameters used in the experiments are given in Table 1.
Every model, including its FANToM variant, is constructed as follows:
SCHOLAR:
SCHOLAR utilizes the common Encoder with the linear Decoder. The softmax of the output of the Encoder, denoted as softmax(enc_out), is used as input to a separate classifier, which predicts document labels.
DVAE:
DVAE employs the common Encoder with the linear Decoder.
ETM:
ETM utilizes the common Encoder with an embedding decoder (Decoder ETM). The word embeddings are initialized with GloVe embeddings.
CTM:
CTM employs the common Encoder with the linear Decoder. The distinction lies in the input to the encoder, which consists of contextualized embeddings rather than Bag of Words (BoW). The contextualized embeddings are sourced from the SBERT model.
ECRTM:
ECRTM employs the common Encoder with the ETM Decoder. The distinction lies in the optimization function, where we use implementation provided by the authors [^1] https://github.com/BobXWu/ECRTM.
All of the aforementioned models are also constructed with their corresponding FANToM variants. In these variants, only the priors are manipulated in the objective function (as mentioned in described in Section LABEL:align-topic-label of the main paper), and a separate author decoder is introduced to learn the topic-author multinomial distribution. The core architecture remains unchanged in all cases.
This also confirms the adaptability of our architecture to seamlessly incorporate any VAE-based topic model. Furthermore, it demonstrates the ability to align the learned topics with expert-assigned labels, and to leverage author information for establishing a meaningful correspondence between topics, authors, and labels.
Table 1: Hyperparameter settings for the experiments.
Hyperparameter Value
batch size 128
$\alpha$ 0.02
$\beta$ 2
Learning Rate 0.001
Max Epochs 100
$p_{keep}$ 0.25
$L^{1}$ $\mathbb{R}^{\text{vocab\_size $\times$ 512}}$
$L^{2}$ $\mathbb{R}^{\text{512 $\times$ total\_topics}}$
$L^{3}$ $\mathbb{R}^{\text{total\_topics $\times$ vocab\_size}}$
$L^{4}$ $\mathbb{R}^{\text{total\_topics $\times$ author\_size}}$
$\delta$ $\mathbb{R}^{\text{total\_words $\times$ 300}}$
$\eta$ $\mathbb{R}^{\text{total\_words $\times$ 300}}$
word embeddings [pennington2014glove, GloVe]
Appendix B Datasets and Preprocessing
For dataset preparation, we use SpaCy [spacy2] for data tokenization. Additionally, we eliminate common stop words, punctuations, as well as high and low-frequency words. High and low-frequency words are determined by excluding words that appear in over 85% of the documents or in fewer than 30 documents. This standardization is applied across all models.
Appendix C Using LLMs as Expert
Not all datasets come with accompanying labels and labeling every corpus can be a challenging task. However, an alternative approach involves using Large Language Models (LLMs) to assign labels to unlabeled documents. In our study, we explore this possibility by using the capabilities of BART [lewis2019bart] for zero-shot text classification. To achieve this, we use a pretrained model from [wolf2019huggingface], which undergoes fine-tuning on Natural Language Inference (NLI) [N18-1101] following the methodology described by [yin2019benchmarking].
This method operates by treating the sequence to be classified as the NLI premise and constructing a hypothesis for each potential label. For example, when determining if a sequence pertains to the ”sport” category, we can formulate a hypothesis such as ”This text is about sports.” By converting the probabilities for entailment and contradiction, we obtain the label probabilities.
We apply this approach to our selected dataset and supply prospective labels from the label set. By comparing the results obtained through this method with the available human labels, we achieve an accuracy of 92%. This validation further supports the practicality and effectiveness of using LLMs as experts for document categorization. Furthermore, we use the LLM labels of the ag news corpus (AGN) and use FANToM(L) to estimate topics. The results demonstrate a mean topic quality of 0.503, surpassing the 0.409 achieved through human labeling. Figure 1 provides an illustration of the topic words estimated by both human and LLM labels. Notably, both sets of topics align well with the corresponding labels.
Figure 1: Comparison between the topic words estimated using human labels (bottom) and LLM labels (top) in the ag news corpus, using FANToM(L) for topic estimation. Both align well with the corresponding labels.
Appendix D Runtime
In this section, we present the runtime of the proposed models compared to existing models. All experiments were conducted on NVIDIA V100 GPUs. The runtimes are presented in Table 2. The results demonstrate that the proposed methods have minimal impact on the overall runtime. It is important to note that only the model training time is considered in the runtime analysis, while data preparation time is excluded. The incorporation of labels and authors introduces a data preparation threshold, but it remains constant and does not scale with the model’s runtime.
Appendix E Additional Evaluation
Apart from the main paper’s evaluation, we provide supplementary assessments. Figure 3 showcases extended qualitative analysis by revealing all the topics acquired by Labeled LDA, existing NTMs, and our proposed FANToM on the 20ng dataset. This illustration highlights the concordance with labels and also showcases the esoteric topics uncovered by our method. It further underscores the suboptimal quality of topics obtained through statistical models.
In addition to qualitative findings, we introduce supplementary quantitative results. Table 4 presents Top-3 and Top-5 accuracy scores of the FANToM(L) model validates the alignment of topics with labels. Table 3 presents F1 scores, affirming the successful alignment. Tables 5 and 6 delve into topic coherence and diversity in comparison to baselines, thus demonstrating the superior quality of topics learned by our model.
Datasets ours (FANToM) baselines
ETM DVAE ETM DVAE
20NG $119.29\pm 1.34$ $119.36\pm 1.46$ $115.84\pm 1.76$ $119.46\pm 1.36$
AGN $157.78\pm 2.02$ $147.48\pm 2.47$ $148.05\pm 3.06$ $143.25\pm 3.54$
DB-14 $155.27\pm 4.03$ $150.84\pm 2.05$ $147.16\pm 2.99$ $144.16\pm 2.79$
arxiv $620.48\pm 4.05$ $605.06\pm 5.88$ $602.99\pm 4.59$ $595.23\pm 6.82$
arxiv $601.23\pm 1.28$ $590.71\pm 8.53$
arxiv∗∗ $656.46\pm 0.69$ $601.85\pm 9.88$
Table 2: This table presents a comparison of the runtime for the proposed FANToM model against baseline models. The upper part of the table displays the runtime values for FANToM(L) when only labels are incorporated, while the bottom part shows the runtime values for FANToM(A) (*) and FANToM (**) respectively. All runtimes are measured in seconds. The results indicate that incorporating the proposed methods does not have a significant impact on the runtime of the models.
Datasets FANToM(L)
%MacroF1 %MicroF1
20NG $0.523\pm 0.02$ $0.541\pm 0.01$
AGN $0.659\pm 0.03$ $0.655\pm 0.03$
DB-14 $0.511\pm 0.00$ $0.532\pm 0.01$
arxiv $0.707\pm 0.03$ $0.725\pm 0.04$
Table 3: F1 scores of the FANToM(L) model validates the alignment of topics with labels.
Datasets FANToM(L)
Top-3 Top-5
20NG $0.696\pm 0.02$ $0.798\pm 0.01$
AGN $0.892\pm 0.02$ $1.000\pm 0.00$
DB-14 $0.726\pm 0.01$ $0.832\pm 0.02$
arxiv $0.852\pm 0.02$ $0.918\pm 0.01$
Table 4: Top-3 and Top-5 accuracy scores of the FANToM(L) model validates the alignment of topics with labels.
Figure 2: Cosine Similarity Matrix illustrating the relationships between prominent ML authors, determined by analyzing their associated topic vectors through the FANToM(A) model. High similarity scores indicate shared research areas among authors.
Figure 3: This figure compares FANToM(L) (top), Labeled LDA (middle), and Neural Topic Models (bottom) on the 20NG dataset. FANToM(L) exhibits aligned and esoteric topics consistent with human intentions, while Labeled LDA shows repetitive generic words in many topics (e.g., ”host,” ”writes,” ”articles”), leading to low diversity. Neural Topic Models do not have association with provided labels and fail to produce certain topics completely, unlike FANToM(L) (e.g., ”mac”, ”graphics”, ”electronics”).
Figure 4: Topics and their top authors from ”stats.ML” category in the arxiv dataset between 2010-2020. The authors are correctly associated with their research area.
Models 20NG AGN DB-14 arxiv
baseline L-LDA $0.198\pm 0.00$ $0.336\pm 0.00$ $0.551\pm 0.00$ $0.269\pm 0.02$
SCHOLAR $0.410\pm 0.01$ $0.372\pm 0.06$ 0.629±0.020.629$\pm$ 0.02$\bullet$0.629 ± 0.02 ∙ $0.364\pm 0.03$
[…]
ETM 0.418±0.030.418$\pm$ 0.03$\bullet$0.418 ± 0.03 ∙ $0.326\pm 0.02$ $0.630\pm 0.01$ $0.281\pm 0.01$
CTM $0.412\pm 0.01$ $0.371\pm 0.01$ 0.632±0.010.632$\pm$ 0.01$\bullet$0.632 ± 0.01 ∙ 0.348±0.020.348$\pm$ 0.02$\bullet$0.348 ± 0.02 ∙
[…]
FANToM-ETM $0.410\pm 0.01$ 0.391±0.020.391$\pm$ 0.02$\bullet$0.391 ± 0.02 ∙ 0.639±0.020.639$\pm$ 0.02$\bullet$0.639 ± 0.02 ∙ 0.310±0.020.310$\pm$ 0.02$\bullet$0.310 ± 0.02 ∙
[…]
Table 5: Comparison of the proposed FANToM(L) (bottom) against the baselines (top) using Topic Coherence metrics. The best results across datasets are highlighted in bold, and best results across corresponding models between baselines and FANToM(L) are marked with $\bullet$. The proposed model generally outperforms the existing baselines.
Models 20NG AGN DB-14 arxiv
baseline L-LDA $0.97\pm 0.00$ $0.95\pm 0.00$ $0.93\pm 0.00$ $0.91\pm 0.01$
SCHOLAR $0.95\pm 0.01$ $1.00\pm 0.00$ $0.95\pm 0.01$ $0.94\pm 0.01$
DVAE $1.00\pm 0.00$ $1.00\pm 0.00$ $0.95\pm 0.01$ $0.94\pm 0.01$
ETM $0.95\pm 0.01$ $0.95\pm 0.00$ $0.96\pm 0.01$ $0.93\pm 0.02$
CTM $1.00\pm 0.00$ $1.00\pm 0.00$ $0.94\pm 0.01$ $0.93\pm 0.02$
ours FANToM-SCHOLAR $0.97\pm 0.01$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
FANToM-DVAE $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
FANToM-ETM $0.97\pm 0.01$ $1.00\pm 0.00$ $0.99\pm 0.00$ $1.00\pm 0.00$
FANToM-CTM $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$ $1.00\pm 0.00$
Table 6: Comparison of the proposed FANToM(L) (bottom) against the baselines (top) using Topic Diversity metrics. The proposed model produces topics with higher diversity.
Appendix F Limitations
One common limitation in the field of topic models is the reliance on automated evaluation measures, including the topic quality measure. While topic coherence, a component of the topic quality measure, has demonstrated a reasonable correlation with human evaluations, it is important to note that qualitative assessment of topics remains subjective, even in human evaluations. Consequently, even with human studies, there is no guarantee of consistent and reproducible results. However, it is worth noting that this challenge is pervasive across the field and not specific to our proposed models. Despite these limitations, our proposed models have consistently performed on par with or outperformed existing state-of-the-art models.
As a second limitation, we rely on labels to align our models. The label quality may vary between datasets, influencing the final results. We have not investigated the quality of the labelings, the human annotations in the datasets, or the LLM-assigned labels. Therefore, the results may be slightly biased depending on the quality of the provided labels.
Appendix G Ethical Statement
In this study, we incorporated authorship information into topic models while ensuring ethical considerations. To acquire the dataset, we obtained it from the openly available [arxiv], which is licensed under the CC0 1.0 Universal (CC0 1.0) Public Domain Dedication, allowing us unrestricted use of the data. We specifically utilized the abstracts and author names of openly available papers. It is important to note that we have mentioned author names as they were presented to us by our model, without making any alterations. We did not have access to any sensitive or restricted information.